An artist is designing a logo for a business in the shape of a circle with an inscribed rectangle.the diameter of circle is 6.5 inches and the area of the rectangle is 15 square inches.Find the dimension of the rectangle?
Found 2 solutions by Boreal, rothauserc:
Answer by Boreal(15235)

The rectangle has to be symmetric with the diameter, so the diagonals have to have diameters of 6.5 in
LW=15; L=15/W
L^2+W^2=6.5^2=42.25
225/W^2+ (W^2)=42.25 by substitution
225+W^4=42.25W^2, multiplying everything by W^2.
W^4-42.25W^2+225=0
4W^4-169W^2+900=0
W^2=(1/8)(169+/- sqrt (28561-14400); sqrt term= 119
W^2=36, W=6 using positive root, L=2.5
(6 inches x 2.5 inches) ANSWER
W^2=6.25, W=2.5 using negative root, L=6, same result.






Answer by rothauserc(4718)

Both diagonals of the rectangle have the same length as the diameter of the circle
:
let l be the length of the rectangle and w be the width, then using the Pythagorean Theorem
:
l^2 + w^2 = 6.5^2 = 42.25
:
l = sqrt(42.25 - w^2)
:
we know the area(A) of the rectangle is 15 square inches and A = l * W, then
:
sqrt(42.25 - w^2) * w = 15
:
square both sides of =
:
(42.25 - w^2) * w^2 = 15
:
42.25w^2 - w^4 = 225
:
w^4 -42.25w^2 +225 = 0
:
let u = w^2, then
:
u^2 -42.25u + 225 = 0
:
(u-36) * (u-6.25) = 0
:
u = 36 or u = 6.25
:
since u = w^2, then w = sqrt(u)
:
w = 6 or 2.5
:
*********************************
we have two possible rectangles
:
l = 15/6 = 2.5, w = 6
:
l = 15/2.5 = 6, w = 2.5
:
********************************
: